OpenStudy (anonymous):

What are the solutions for x in the proportion 5x-6 (over) x = 2x (over)2

OpenStudy (king):

x=3,2

OpenStudy (anonymous):

(5x-6)/x = 2x/2 => (5x - 6)/x = x \[x ^{2} - 5x + 6 = 0 => x= 2 or x = 3\]

OpenStudy (king):

use quad. equation formula

OpenStudy (anonymous):

thank you